# Stability of quasicrystalline ultracold fermions to dipolar interactions

### v2

@@ Abstract
Introduction — Since their revolutionary discovery in aluminum-manganese alloys Shechtman et al. (1984), quasicrystals have been the subject of intense theoretical and experimental research due to their many exotic properties. Quasicrystals exhibit a regular structure despite lacking a canonical periodicity Rabson et al. (1991); Goldman and Kelton (1993). While they lack translational symmetry, they can completely tesselate a two-dimensional surface in an aperiodic fashion because they do possess unexpected rotational symmetries (five-fold, eight-fold, etc.). Quasicrystals can be mathematically understood in terms of higher-dimensional periodic structures projected onto lower-dimensional slices, and have thus ushered in new paradigms for crystallography Mackay (1995); Yamamoto (1996); Steurer and Deloudi (2008, 2009); Wolny et al. (2018), band theory in solids Poon (1992); Kelton (1993); Rotenberg et al. (2004); Grimm and Schreiber (2003), and hidden dimensions Jagannathan (2021). Moreover, quasiperiodic structures can emulate on-site disorder, which makes them ideal platforms to study many-body localization Iyer et al. (2013); Schreiber et al. (2015); Setiawan et al. (2017); Khemani et al. (2017); Zhang and Yao (2018); Macé et al. (2019); Varma and Žnidarič (2019); Abanin et al. (2019); Agrawal et al. (2022) and its interplay with nonergodicity Li et al. (2015, 2016); Hsu et al. (2018); Ghosh et al. (2020); Roy and Sharma (2021); Roy et al. (2023), integrability He et al. (2013); Žnidarič and Ljubotina (2018); Singh et al. (2021); Žnidarič (2021); Thomson and Schirò (2023), and fractality Thiem and Schreiber (2011); Kalugin and Katz (2014); dimensions of wave functions et al. (2016); Macé et al. (2017); Rai et al. (2019, 2020); Jagannathan (2021); Dai et al. (2023).
@@
In recent years, impressive advances in trapping and controlling ultracold atoms have made them an excellent platform where to realize and study quasiperiodicity with extreme precision Lye et al. (2007); Fallani et al. (2007); Singh et al. (2015); Schreiber et al. (2015); Macé et al. (2016); Bordia et al. (2017); Lüschen et al. (2017a, b); Lueschen et al. (2018); Corcovilos and Mittal (2019); Viebahn et al. (2019); Johnstone et al. (2019); Sbroscia et al. (2020); Gautier et al. (2021); Nakajima et al. (2021); Yu et al. (2023). In one dimension (1D), a quasiperiodic potential can be generated by superimposing a primary optical lattice with a detuning lattice at an incommensurate frequency. Upon loading particles in such a quasiperiodic potential, it is possible to quantum simulate and study the dynamics of 1D quasicrystals with high accuracy. A particularly interesting feature exhibited by these systems is the coexistence of localized and extended states in an intermediate phase separating the fully localized to the fully extended phase Lueschen et al. (2018). In the intermediate phase, the different states are separated by a critical energy level termed single-particle mobility edge (SPME) Abrahams et al. (1979); Lee and Ramakrishnan (1985); Boers et al. (2007); Ganeshan et al. (2015); Li et al. (2017); Lueschen et al. (2018). This coexistence manifests itself due to long-range tunneling terms that appear in the continuum description of the 1D ultracold quantum simulator. In fact, the SPME is absent in the deep lattice limit, where an effective Aubry-André model with self-duality is obtained Aubry and André (1980); Fallani et al. (2007); Roati et al. (2008); Modugno (2009). Furthermore, a rigorous SPME does not appear in 1D systems with Anderson-type (uncorrelated) disorder Lueschen et al. (2018).
@@
with $\mathcal{N}_{c}=\int_{c}\mathrm{d}x\>\rho(x,t)$ and $c$ referring to the initially populated central region (e.g. for $N=4$, $c=[-4.5,3.5]$). These two observables in the long-time limit help us to determine the many-body phase of the system. A finite imbalance $\mathcal{I}$ indicates the persistence of the initial charge density wave pattern imposed by the superlattice and is a proxy for the presence of localized states. A finite (and growing) value of the edge density $\mathcal{D}$ reveals instead the presence of extended states. Therefore, in the localized phase $\mathcal{I}>0$ and $\mathcal{D}=0$, while in the extended phase $\mathcal{I}=0$ and $\mathcal{D}>0$. Localized and extended states can also coexist at different energies in an intermediate phase, where both values of $\mathcal{I}$ and $\mathcal{D}$ are nonzero and a SPME is present. While the SPME should occur at arbitrarily small nonzero $\mathcal{I}$ and $\mathcal{D}$, numerical simulations introduce small imprecisions. Thus small but finite thresholds $t_{\mathcal{I}}$ and $t_{\mathcal{D}}$ for the imbalance and the edge density have to be introduced. Throughout our analysis, we will empirically set them to be $t_{\mathcal{I}}=0.1$ and $t_{\mathcal{D}}=0.03$  [^1] We found empirically that these threshold values minimize numerical deviations from the experimentally observed phase diagram..
@@
Methods — To simulate the full many-body interacting system, we employ the MultiConfigurational Time-Dependent Hartree method for indistinguishable particles Streltsov et al. (2006, 2007); Alon et al. (2007, 2008), implemented in the MCTDH-X software Lode (2016); Fasshauer and Lode (2016); Lin et al. (2020a); Lode et al. (2020, 2024). With it, we solve the many-body Schrödinger equation directly for the continuum system consisting of kinetic energy, interparticle interactions, and optical lattice. MCTDH-X has been widely used to study ground-state and dynamical properties of long-range interacting systems Fischer et al. (2015); Molignini et al. (2018); Chatterjee and Lode (2018); Lin et al. (2019); Chatterjee et al. (2019, 2020); Lin et al. (2020b, 2021); Molignini et al. (2022); Hughes et al. (2023); Bilinskaya et al. (2024); Molignini and Chakrabarti (2024a, b). MCTDH-X relies on a time-dependent variational optimization procedure in which the many-body wavefunction is decomposed into an adaptive basis set of $M$ time-dependent single-particle functions called orbitals. Due to the Pauli exclusion principle, to correctly describe a set of $N$ fermions, we require $M\geq N+1$ orbitals  [^2] Strictly speaking, $M=N$ would suffice but this case is often degenerate.. For strongly interacting systems, a larger number of orbitals might be required to capture many-body correlation effects. In this work, we have verified that including orbitals beyond $M=N+4$ leads to a negligible population for the time scales probed by our dynamics sup .
@@
Results for noninteracting fermions — We begin by describing the dynamics for noninteracting particles. Fig. 1(a) shows the noninteracting phase diagrams for $N=4$ fermions with $M=8$ orbitals, constructed by calculating the imbalance and expansion observables as the depth of the primary and detuning lattices are varied. We use parameter ranges compatible with experimental regimes, i.e. $V_{p}\in[3E_{r},8E_{r}]$ and $V_{d}\in[0,1.2E_{r}]$. The imbalance is calculated by averaging over the last 50 time steps at the end of each time evolution. The edge density is calculated at a target time $\bar{t}$, well before the particles can reach the boundaries of the system. Typical target times are $\bar{t}=0.45$ ms to $\bar{t}=1.78$ ms  [^3] This measurement timescale is slightly different than the experimental one, where expansion measurements are calculated at longer times. This is necessary to avoid simulating very large spatial grids that are computationally unattainable. Nevertheless, we have verified that the value of the target time $\bar{t}$ only slightly impacts the precise location of the phase boundary between the intermediate and localized phase, and that the qualitative behavior is the same for different values of $\bar{t}$.. From Fig. 1(a), we can clearly distinguish the three expected phases, which have the same structure observed in prior studies Li et al. (2017); Lueschen et al. (2018). At low values of $V_{d}$, the system is always in the extended phase (purple region). Increasing the value of $V_{d}$ at low to moderate values of $V_{p}$ pushes the system into an intermediate region where extended and localized states coexist (magenta region). Further increasing $V_{d}$ or $V_{p}$ eventually localizes all the states (yellow region).
@@
Figure 1: Behavior of the noninteracting system: (a) Phase diagram for $N=4$ fermions with $M=8$ orbitals in a 1D quasiperiodic potential. For the localized phase, the time-averaged imbalance $\mathcal{I}\right>$ is plotted in a continuous color scheme. (b)-(d) 2-RDM at $t=0.89$ ms for (a) localized phase ($V_{p}=8.0E_{r}$, $V_{d}=1.2E_{r}$), (b) intermediate phase ($V_{p}=6.0E_{r}$, $V_{d}=0.4E_{r}$), and (c) extended phase ($V_{p}=3.0E_{r}$, $V_{d}=0E_{r}$).
@@
The three phase of the quasicrystal system can also be distinguished by the dynamics of their many-body correlations encoded in the 2-RDM, shown for selected times in Fig. 1(b)-(d) (see supplementary materials for more time snapshots sup ). The 2-RDM in the localized phase [Fig. 1(b)] does not change at all with time and merely shows the expected exchange hole at $x=x^{\prime}$ with strong many-body correlation between the sites of the initial superlattice. On the contrary, the extended phase [Fig. 1(d)] exhibits a rapid spread of correlations at very short times already. Between the two extremes of localization and expansion lies the intermediate phase [Fig. 1(c)]. The correlation pattern here reveals an almost immediate exchange correlation between neighboring sites within the populated area of the initial superlattice, and a small spread to sites just outside of it. However, the intensity of the 2-RDM is not uniform: correlations between sites in the initial superlattice structure still dominate throughout the time evolution even for longer times sup . These results indicate that the intermediate phase is characterized by two types of correlations showcasing the coexistence of localized and extended states. Localized states mediate static, short-range correlations between the initially populated sites, while extended states mediate long-range correlations spanning many sites in the quasiperiodic lattice.
@@
Figure 2: Phase diagram for increasing DDI strength $W$: (a) $W=0.02$$E_{r}$, (b) $W=0.2$$E_{r}$, (c) $W=2.0$$E_{r}$. For the localized phase, the time-averaged imbalance parameter $\mathcal{I}\right>$ is plotted as a continuous color scheme. The cyan symbols show the parameters used in Fig. 3. The green triangles indicate the points for which the density dynamics is plotted in panels (d)-(f) (note that for (d)-(e) the maximal density value shown is purposefully small to highlight boundary fluctuations).
Fig. 2 shows how the noninteracting phase diagram changes when interactions are progressively increased [panels (a) to (c)]. At small interaction values, the intermediate phase only grows slightly while the imbalance in the localized phase decreases. As the interactions become stronger, though, the phase diagram changes drastically and in a non-monotonic fashion. For $W=0.2E_{r}$, the intermediate phase grows sizeably at the expense of most of the localized phase and part of the extended phase. Therefore, DDI stabilize the coexistence of extended and localized states in the intermediate phase. At even stronger interactions, $W=2.0E_{r}$, the intermediate phase dominates the probed parameter region. Note that in Fig. 2(c) $V_{p}$ is plotted to larger values up to $12E_{r}$. The purely extended region is instead reduced to a tiny pool around $V_{p}=3E_{r}$, $V_{d}=0E_{r}$. Most strikingly, however, a new localized region emerges at low values of $V_{d}$. Note that this region was initially part of the extended phase in the noninteracting case. The localized phases at low and high $V_{d}$ are divided by a strip of intermediate phase which persists for very large values of $V_{p}$ (as large as $V_{p}=16.0E_{r}$ – not shown here), reminiscent of the Arnold tongues found in parametrically driven systems McLachlan (1951); Magnus and Winkler (1966); Molignini et al. (2018). In this lobe, the particles undergo density oscillations around their initial pinned configuration, indicating a potential resonance between DDI and quasiperiodic potential [see Fig. 2(f)]. While both regions above and below the intermediate phase lobe show strong localization, their behavior is different. Above the lobe (high $V_{d}$, Fig. 2(d)), the particles do not move at all from their initial configuration. Below the lobe (low $V_{d}$, Fig. 2(e)), the particle density periodically leaks back and forth from the initial boundaries while retaining a strong localization inside them, indicating a dynamically self-bound state (DSBS). Such states were recently discovered in bosonic lattice models with attractive DDI Barbiero et al. (2015); Li et al. (2020, 2021a, 2021b); Korbmacher et al. (2023). Here, we reveal that DSBS can also exist for fermionic gases with repulsive DDI and away from the lattice limit $V_{d}\to\infty$.
@@
Figure 3: Dynamics of dipolar fermions with increasing DDI strength $W$ in a quasiperiodic lattice. The interaction strength is (a)-(c) $W=0.02E_{r}$, (d)-(f) $W=0.2E_{r}$, (g)-(i) $W=2.0E_{r}$. The potential depths are (a), (d), (g) $V_{p}=3.0E_{r}$, $V_{d}=0.0E_{r}$ (extended phase) (b), (e), (h) $V_{p}=8.0E_{r}$, $V_{d}=1.2E_{r}$ (localized phase), (c) $V_{p}=7.0E_{r}$, $V_{d}=0.2E_{r}$, (f) $V_{p}=6.0E_{r}$, $V_{d}=0.4E_{r}$, (i) $V_{p}=5.0E_{r}$, $V_{d}=0.6E_{r}$ (intermediate phase). The red symbols indicate the location of the parameters in Fig. 2. The axes in (g) apply to all panels.
@@
The increase in localization due to DDI interference is also observed in the 2-RDM. This is shown in Fig. 4 for the three different phases and at three increasing values of $W$. The behavior for weak DDI [$W=0.02E_{r}$, Fig. 4(a)-(c)] is similar to what already observed in the noninteracting case. However, when DDI are increased [$W=0.2E_{r}$ in Fig. 4(d)-(f) and $W=2.0E_{r}$ in Fig. 4(g)-(i)], the correlation spread becomes progressively stiffer for all phases. This leads to a complete freeze of correlations in the localized phase [Fig. 4(g)], while in the intermediate phase correlations are able to spread only partially  [^4] We attribute this unidirectional spread to pinning by the given choice of the quasiperiodic potential. Thus, the location of the pinning might change with a different choice of the wave vector ratio and detuning phase.. Even for the extended phase, the pinning effect is visible, with correlations spreading more easily towards the right than the left. This observation agrees with the uneven density spread observed in Fig. 3(g).
Figure 4: 2-RDM in the dynamics of fermions with DDI. The snapshots are taken at $t=0.89$ ms and for increasing DDI: (a)-(c) $W=0.02E_{r}$, (d)-(f) $W=0.2$$E_{r}$, (g)-(i) $W=2.0E_{r}$. The parameters are (a), (d), (g) $V_{p}=8.0E_{r}$, $V_{d}=1.2E_{r}$ (localized phase), (b), (e), (h) $V_{p}=6.0E_{r}$, $V_{d}=0.4E_{r}$ (intermediate phase), (c), (f), (i) $V_{p}=3.0E_{r}$, $V_{d}=0E_{r}$ (extended phase). The axes in (g) apply to all panels.
@@
Discussion — There are several mechanisms at play that explain the overall shape of the phase diagram at strong DDI. The localized state appearing at large $V_{p}$ (lattice limit) in the clean limit $V_{d}\to 0$ is a DSBS stabilized by DDI. The $1/r^{3}$ tails of the DDI cause an interference that locks particles within a certain distance, as demonstrated in similar bosonic systems Barbiero et al. (2015); Li et al. (2020, 2021a, 2021b); Korbmacher et al. (2023), and is likely connected with Hilbert space fragmentation Khemani et al. (2020); Herviou et al. (2021); Cheng (2023). Away from the clean limit ($V_{d}\neq 0$, but $V_{d}\ll V_{p}$), correlated disorder can be regarded as a perturbation on top of the DSBS. In this setting, it is known that the DDI give rise to effective correlated hopping at a distance, i.e. between the edges of the localized clusters Levitov (1989, 1990); Burin and Kagan (1994); Burin (2015a, b, 2017); Nosov et al. (2019), which explains the resonant-like behavior at the edges at their progressive delocalization. Increasing the disorder further eventually pushes the system again into localization. In this limit, the system approaches an Aubry-André model, whose localized phase is known to be stable against DDI Gopalakrishnan (2017); Nag and Garg (2019); Deng et al. (2019).
@@
Acknowledgments — We acknowledge computation time on the ETH Zurich Euler cluster and at the High-Performance Computing Center Stuttgart (HLRS). We thank O. Alon, E. Bergholtz, and R. Chitra, E. Fasshauer, and R. Lin for useful discussions. This work is supported by the Swedish Research Council (2018-00313) and Knut and Alice Wallenberg Foundation (KAW) via the project Dynamic Quantum Matter (2019.0068).
Supplemental Material for:

### v3

@@ Abstract
Introduction — Since their revolutionary discovery in aluminum-manganese alloys Shechtman et al. (1984), quasicrystals have been the subject of intense theoretical and experimental research due to their many exotic properties. Quasicrystals exhibit a regular structure despite lacking a canonical periodicity Rabson et al. (1991); Goldman and Kelton (1993). While they lack translational symmetry, they can completely tesselate a two-dimensional surface in an aperiodic fashion because they do possess unexpected rotational symmetries (five-fold, eight-fold, etc.). Quasicrystals can be mathematically understood in terms of higher-dimensional periodic structures projected onto lower-dimensional slices, and have thus ushered in new paradigms for crystallography Mackay (1995); Yamamoto (1996); Steurer and Deloudi (2008, 2009); Wolny et al. (2018), band theory in solids Poon (1992); Kelton (1993); Rotenberg et al. (2004); Grimm and Schreiber (2003), and hidden dimensions Jagannathan (2021). Moreover, quasiperiodic structures can emulate on-site disorder, which makes them ideal platforms to study many-body localization Iyer et al. (2013); Schreiber et al. (2015); Setiawan et al. (2017); Khemani et al. (2017); Zhang and Yao (2018); Macé et al. (2019); Varma and Žnidarič (2019); Abanin et al. (2019); Agrawal et al. (2022) and its interplay with nonergodicity Li et al. (2015, 2016); Hsu et al. (2018); Ghosh et al. (2020); Roy and Sharma (2021); Roy et al. (2023), integrability He et al. (2013); Žnidarič and Ljubotina (2018); Singh et al. (2021); Žnidarič (2021); Thomson and Schirò (2023), and fractality Thiem and Schreiber (2011); Kalugin and Katz (2014); Macé et al. (2016a, 2017); Rai et al. (2019, 2020); Jagannathan (2021); Dai et al. (2023).
@@
In recent years, impressive advances in trapping and controlling ultracold atoms have made them an excellent platform where to realize and study quasiperiodicity with extreme precision Lye et al. (2007); Fallani et al. (2007); Singh et al. (2015); Schreiber et al. (2015); Macé et al. (2016b); Bordia et al. (2017); Lüschen et al. (2017a, b); Lueschen et al. (2018); Corcovilos and Mittal (2019); Viebahn et al. (2019); Johnstone et al. (2019); Sbroscia et al. (2020); Gautier et al. (2021); Nakajima et al. (2021); Yu et al. (2024). In one dimension (1D), a quasiperiodic potential can be generated by superimposing a primary optical lattice with a detuning lattice at an incommensurate frequency. Upon loading particles in such a quasiperiodic potential, it is possible to quantum simulate and study the dynamics of 1D quasicrystals with high accuracy. A particularly interesting feature exhibited by these systems is the coexistence of localized and extended states in an intermediate phase separating the fully localized to the fully extended phase Lueschen et al. (2018). In the intermediate phase, the different states are separated by a critical energy level termed single-particle mobility edge (SPME) Abrahams et al. (1979); Lee and Ramakrishnan (1985); Boers et al. (2007); Ganeshan et al. (2015); Li et al. (2017); Lueschen et al. (2018). This coexistence manifests itself due to long-range tunneling terms that appear in the continuum description of the 1D ultracold quantum simulator. In fact, the SPME is absent in the deep lattice limit, where an effective Aubry-André model with self-duality is obtained Aubry and André (1980); Fallani et al. (2007); Roati et al. (2008); Modugno (2009). Furthermore, a rigorous SPME does not appear in 1D systems with Anderson-type (uncorrelated) disorder Lueschen et al. (2018).
@@
with $\mathcal{N}_{c}=\int_{c}\mathrm{d}x\>\rho(x,t)$ and $c$ referring to the initially populated central region (e.g. for $N=4$, c=[4.5,3.5]c=[-4.5,3.5]). These two observables in the long-time limit help us to determine the many-body phase of the system. A finite imbalance $\mathcal{I}$ indicates the persistence of the initial charge density wave pattern imposed by the superlattice and is a proxy for the presence of localized states. A finite (and growing) value of the edge density $\mathcal{D}$ reveals instead the presence of extended states. Therefore, in the localized phase $\mathcal{I}>0$ and $\mathcal{D}=0$, while in the extended phase $\mathcal{I}=0$ and $\mathcal{D}>0$. Localized and extended states can also coexist at different energies in an intermediate phase, where both values of $\mathcal{I}$ and $\mathcal{D}$ are nonzero and a SPME is present. While the SPME should occur at arbitrarily small nonzero $\mathcal{I}$ and $\mathcal{D}$, numerical simulations introduce small imprecisions. Thus small but finite thresholds $t_{\mathcal{I}}$ and $t_{\mathcal{D}}$ for the imbalance and the edge density have to be introduced. Throughout our analysis, we will empirically set them to be $t_{\mathcal{I}}=0.1$ and $t_{\mathcal{D}}=0.03$  [^1] We found empirically that these threshold values minimize numerical deviations from the experimentally observed phase diagram..
@@
Methods — To simulate the full many-body interacting system, we employ the MultiConfigurational Time-Dependent Hartree method for indistinguishable particles Streltsov et al. (2006, 2007); Alon et al. (2007, 2008), implemented in the MCTDH-X software Lode (2016); Fasshauer and Lode (2016); Lin et al. (2020a); Lode et al. (2020, 2024). With it, we solve the many-body Schrödinger equation directly for the continuum system consisting of kinetic energy, interparticle interactions, and optical lattice. MCTDH-X has been widely used to study ground-state and dynamical properties of long-range interacting systems Fischer et al. (2015); Molignini et al. (2018); Chatterjee and Lode (2018); Lin et al. (2019); Chatterjee et al. (2019, 2020); Lin et al. (2020b, 2021); Molignini et al. (2022); Hughes et al. (2023); Bilinskaya et al. (2024); Molignini and Chakrabarti (2025, 2024). MCTDH-X relies on a time-dependent variational optimization procedure in which the many-body wavefunction is decomposed into an adaptive basis set of $M$ time-dependent single-particle functions called orbitals. Due to the Pauli exclusion principle, to correctly describe a set of $N$ fermions, we require $M\geq N+1$ orbitals  [^2] Strictly speaking, $M=N$ would suffice but this case is often degenerate.. For strongly interacting systems, a larger number of orbitals might be required to capture many-body correlation effects. In this work, we have verified that including orbitals beyond $M=N+4$ leads to a negligible population for the time scales probed by our dynamics sup .
@@
Results for noninteracting fermions — We begin by describing the dynamics for noninteracting particles. Fig. 1(a) shows the noninteracting phase diagrams for $N=4$ fermions with $M=9$ orbitals, constructed by calculating the imbalance and expansion observables as the depth of the primary and detuning lattices are varied. We use parameter ranges compatible with experimental regimes, i.e. $V_{p}\in[3E_{r},8E_{r}]$ and $V_{d}\in[0,1.2E_{r}]$. The imbalance is calculated by averaging over the last 50 time steps at the end of each time evolution, with maximal simulation times reaching $\approx 45$ ms. The edge density is typically calculated at a target time $\tilde{t}$, well before the particles can reach the boundaries of the system. Typical target times are $\tilde{t}=1.8$ ms to $\tilde{t}=7.2$ ms  [^3] This measurement timescale is slightly different than the experimental one, where expansion measurements are calculated at longer times. This is necessary to avoid simulating very large spatial grids that are computationally unattainable. Nevertheless, we have verified that the value of the target time $\bar{t}$ only slightly impacts the precise location of the phase boundary between the intermediate and localized phase, and that the qualitative behavior is the same for different values of $\bar{t}$.. From Fig. 1(a), we can clearly distinguish the three expected phases, which have the same structure observed in prior studies Li et al. (2017); Lueschen et al. (2018). At low values of $V_{d}$, the system is always in the extended phase (purple region). Increasing the value of $V_{d}$ at low to moderate values of $V_{p}$ pushes the system into an intermediate region where extended and localized states coexist (magenta region). Further increasing $V_{d}$ or $V_{p}$ eventually localizes all the states (yellow region).
@@
Figure 1: Behavior of the noninteracting system: (a) Phase diagram for $N=4$ fermions with $M=9$ orbitals in a 1D quasiperiodic potential. For the localized phase, the time-averaged imbalance $\left$$\mathcal{I}\right$> is plotted in a continuous color scheme. (b)-(d) 2-RDM at $t=20\bar{t}=3.7$ ms for (b) localized phase ($V_{p}=16.0E_{r}$, $V_{d}=2.4E_{r}$), (c) intermediate phase ($V_{p}=5.0E_{r}$, $V_{d}=0.4E_{r}$), and (d) extended phase ($V_{p}=3.0E_{r}$, $V_{d}=0E_{r}$).
@@
The three phases of the quasicrystal system can also be distinguished by the dynamics of their many-body correlations encoded in the 2-RDM, shown for selected times in Fig. 1(b)-(d) (see supplementary materials for more time snapshots sup ). The 2-RDM in the localized phase [Fig. 1(b)] does not change at all with time and merely shows the expected exchange hole at $x=x^{\prime}$ with strong many-body correlation between the sites of the initial superlattice. On the contrary, the extended phase [Fig. 1(d)] exhibits a rapid spread of correlations at very short times already. Between the two extremes of localization and expansion lies the intermediate phase [Fig. 1(c)]. The correlation pattern here reveals an almost immediate exchange correlation between neighboring sites within the populated area of the initial superlattice, and a relatively small spread to sites outside of it. However, the intensity of the 2-RDM is not uniform: correlations between sites in the initial superlattice structure still dominate throughout the time evolution even for longer times sup . These results indicate that the intermediate phase is characterized by two types of correlations showcasing the coexistence of localized and extended states. Localized states mediate static, short-range correlations between the initially populated sites, while extended states mediate long-range correlations spanning many sites in the quasiperiodic lattice.
@@
[…]
Fig. 2 shows how the noninteracting phase diagram changes when interactions are progressively increased [panels (a) to (c)]. At small interaction values, the intermediate phase only grows slightly while the imbalance in the localized phase decreases. As the interactions become stronger, though, the phase diagram changes drastically and in a non-monotonic fashion. For $W=0.2E_{r}$, the intermediate phase grows sizeably at the expense of most of the localized phase and part of the extended phase. Therefore, DDI stabilize the coexistence of extended and localized states in the intermediate phase. At even stronger interactions, $W=2.0E_{r}$, the intermediate phase dominates the probed parameter region. Note that in Fig. 2(c) $V_{p}$ is plotted to larger values up to $12E_{r}$. The purely extended region is instead reduced to a tiny pool around $V_{p}=3E_{r}$, $V_{d}=0E_{r}$.
Most strikingly, however, a new localized region emerges at low values of $V_{d}$. Note that this region was initially part of the extended phase in the noninteracting case. The localized phases at low and high $V_{d}$ are divided by a strip of intermediate phase which persists for very large values of $V_{p}$ (as large as $V_{p}=16.0E_{r}$ – not shown in the phase diagram), reminiscent of the Arnold tongues found in parametrically driven systems McLachlan (1951); Magnus and Winkler (1966); Molignini et al. (2018). In this lobe, the particles undergo density oscillations around their initial pinned configuration, indicating a potential resonance between DDI and quasiperiodic potential [see Fig. 2(d) and Fig. 3(b)-(d)]. While both regions above and below the intermediate phase lobe show strong localization, their behavior is different. Above the lobe (high $V_{d}$, Fig. 3(e)-(f)), the particles do not move at all from their initial configuration. Below the lobe (low $V_{d}$, Fig. 3(a)), the particle density periodically leaks back and forth from the initial boundaries while retaining a strong localization inside them, indicating a dynamically self-bound state (DSBS). Such states were recently discovered in bosonic lattice models with attractive DDI Barbiero et al. (2015); Li et al. (2020, 2021a, 2021b); Korbmacher et al. (2023). Here, we reveal that DSBS can also exist for fermionic gases with repulsive DDI and away from the lattice limit $V_{d}\to\infty$.
Figure 3: Two types of localization at strong interactions. Density dynamics at $V_{p}=8E_{r}$ and $W=2.0E_{r}$ for increasing values of $V_{d}$ between 0 and $1.0E_{r}$, illustrating how correlated disorder impacts the DSBS. The scale is zoomed in to increase resolution.
@@
Figure 4: Dynamics of dipolar fermions with increasing DDI strength $W$ in a quasiperiodic lattice. The interaction strength is (a)-(c) $W=0.02E_{r}$, (d)-(f) $W=0.2E_{r}$, (g)-(i) $W=2.0E_{r}$. The potential depths are (a), (d), (g) $V_{p}=3.0E_{r}$, $V_{d}=0.0E_{r}$ (extended phase) (b), (e), (h) $V_{p}=8.0E_{r}$, $V_{d}=1.2E_{r}$ (localized phase), (c) $V_{p}=7.0E_{r}$, $V_{d}=0.2E_{r}$, (f) $V_{p}=6.0E_{r}$, $V_{d}=0.4E_{r}$, (i) $V_{p}=5.0E_{r}$, $V_{d}=0.6E_{r}$ (intermediate phase). The axes in (g) apply to all panels.
@@
The increase in localization due to DDI interference is also observed in the 2-RDM. This is shown in Fig. 5 for the three different phases and at three increasing values of $W$. The behavior for weak DDI $W=0.02E_{r}$, Fig. 5(a)-(c)] is similar to what already observed in the noninteracting case. However, when DDI are increased $W=0.2E_{r}$ in Fig. 5(d)-(f) and $W=2.0E_{r}$ in Fig. 5(g)-(i)], the correlation spread becomes progressively stiffer for all phases. This leads to a complete freeze of correlations in the localized phase [Fig. 5(g)], while in the intermediate phase correlations are able to spread only partially  [^4] We attribute this unidirectional spread to pinning by the given choice of the quasiperiodic potential. Thus, the location of the pinning might change with a different choice of the wave vector ratio and detuning phase.. Even for the extended phase, the pinning effect is visible, with correlations spreading more easily towards the right than the left. This observation agrees with the uneven density spread observed in Fig. 4(g).
[…]
@@
Discussion — There are several mechanisms at play that explain the overall shape of the phase diagram at strong DDI. The localized state appearing at large $V_{p}$ (lattice limit) in the clean limit $V_{d}\to 0$ is a DSBS stabilized by DDI, as evinced by the small density oscillations at the boundaries suggesting dynamical self-confinement [see Fig. 3(a)]. The $1/r^{3}$ tails of the DDI cause an interference that locks particles within a certain distance, as demonstrated in similar bosonic systems Barbiero et al. (2015); Li et al. (2020, 2021a, 2021b); Korbmacher et al. (2023), and is likely connected with Hilbert space fragmentation Khemani et al. (2020); Herviou et al. (2021); Cheng (2023). Away from the clean limit ($V_{d}\neq 0$, but $V_{d}\ll V_{p}$), correlated disorder can be regarded as a perturbation on top of the DSBS. In this setting, it is known that the DDI give rise to effective correlated hopping at a distance, i.e. between the edges of the localized clusters Levitov (1989, 1990); Burin and Kagan (1994); Burin (2015a, b, 2017); Nosov et al. (2019), which explains the resonant-like behavior at the edges at their progressive delocalization [Fig. 3(b)-(d)]. Increasing the disorder further eventually pushes the system again into localization [Fig. 3(f)]. In this limit, the system approaches an Aubry-André model, whose localized phase is known to be stable against DDI Gopalakrishnan (2017); Nag and Garg (2019); Deng et al. (2019).
@@
Acknowledgments — We acknowledge computation time on the ETH Zurich Euler cluster, at the High-Performance Computing Center Stuttgart (HLRS), and on the SUNRISE cluster of Stockholm University. We thank O. Alon, E. Bergholtz, R. Chitra, E. Fasshauer, I. Khaymovich, and R. Lin for useful discussions. This work is supported by the Swedish Research Council (grants 2018-00313 and 2024-05213) and Knut and Alice Wallenberg Foundation (KAW) via the project Dynamic Quantum Matter (2019.0068).
Supplemental Material for